OpenStudy (anonymous):

Plzz help!! Which of the following statements best describes the graph of -5x + 2y = 1?

OpenStudy (anonymous):

A. It is a curve joining the points (-5, 2), (2, 3), and (4, 1). B. It is a curve joining the points (-1, -3), (-1, -3), and (1, 5). C. It is a straight line joining the points (1, 3), (3, 8), and (-3, -7). D. It is a straight line joining the points (4, -3), (-1, 2), and (-4, 5).

OpenStudy (zale101):

\(-5x + 2y = 1\) can this function be expressed in a slope form \(y=mx+b\) ?

OpenStudy (zale101):

@Calvin3

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Yes it can

OpenStudy (anonymous):

If you inverse -5x (which adds 5x to both sides) to get 2y alone and divide both sides by 2 to get y completely alone, The expression can be explained in a y=mx+b as y=2/5x + 1/2

OpenStudy (anonymous):

5/2*

OpenStudy (anonymous):

so then the answer would be A.?

OpenStudy (anonymous):

When the slope-intercept form is implemented, more often then not, it will not be curved.

OpenStudy (anonymous):

so then it's not A or B

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so that leaves C and D could you help me find out which one it is

OpenStudy (anonymous):

I tried plotting them and C came out right, so I would say C

OpenStudy (anonymous):

Thx for all your help :)